One number is 5 greater than another number. If the smaller number is added to twice the larger number, the answer is 28. Find the two numbers
Answer by sachi(548)

let the small no=x
the larger no=x+5
now x+2(x+5)=28
or 3x=28-10=18
or x=6 & the larger no=11
ans